### v1

@@ Abstract
Individual behavioural heterogeneity is key to understanding how natural populations adapt flexibly to environmental changes. In ant species like Aphaenogaster senilis, two key roles in collective foraging are commonly recognised: scouts, who discover food patches, and recruits, who exploit these patches and transport food back to the nest. These roles involve distinct movement patterns and exploratory behaviours. In this chapter, we develop a correlated random walk model on the bounded honeycomb lattice to interpret and replicate empirical observations of foraging ants in an enclosed arena with honeycomb tiling. We do so by extending the theory of first-passage processes for $\mathcal{N}$ random walkers when individuals belong to a heterogeneous population. We apply this theory to examine how individual behavioural heterogeneity in ants affects collective foraging efficiency, focusing on first-passage time statistics for nest-to-patch and patch-to-patch movements. With the combined use of the mathematical model and the controlled experimental setup we evaluate (i) the impact of distinct movement strategies by scouts and recruits on finding food items and (ii) whether ants practice strict central place foraging or utilise previously discovered patches as starting points for further exploration.
@@ 1 Introduction
Social groups are rarely uniform or cohesive units; instead, individuals within these groups often exhibit consistent behavioural differences, commonly referred to as “personalities” Jolles2020 . Understanding individual heterogeneity is essential for grasping the variability in life histories and population dynamics chesson1991 . Moreover, it significantly influences the functionality of key ecological responses. Individual heterogeneity is also evident in eusocial species, such as ants and termites, where different roles are associated with different tasks antecologybook , e.g., typically one female and several males are reproductively active, while nonbreeding individuals care for the young, protect the colony, and provide resources. A key debate in this regard is whether these roles represent specialised task forces designed to address different needs or if individuals switch between tasks to adapt to inherent environmental uncertainties Gordon2016 .
In the context of foraging processes, ants are well known to have evolved various strategies for efficient localisation and exploitation of food resources antecologybook . Two common strategies are group and mass recruitment. In group recruitment, a scout discovers a food source and returns to the nest to recruit a small number of nestmates by leaving a chemical trail that only the scout can trace. This strategy, where nestmates are followers, suits smaller or scattered food sources. On the other hand, mass recruitment, where the scout lays a strong pheromone trail that attracts many nestmates, is faster and involves more ants, ideal for large or abundant food sources. Key determinants to select the appropriate strategy are the environmental conditions and the type of food available. For instance, some species primarily rely on individual foraging but switch to group or mass recruitment for richer or larger food sources, allowing ant colonies to optimise resource acquisition across different scenarios. The number of scouts and recruits involved in food collection is also flexible, varying dynamically as the food is being collected and with the amount and predictability of resources. This adaptability raises intriguing questions about the role of these distinct behaviours, how colonies adjust their numbers, and how these adjustments impact collective success.
@@
A. senilis has been selected because of its known group recruitment foraging strategies whereby scouts and recruits have distinct roles campos2014 ; martin2024 . Recent studies on a Y-maze arena have identified significant disparities in the movement patterns of these two types of ants pol2024liquid , with quantifiable differences in the probabilities of moving left, right, or backward at each step. Scouts exhibit less sinuous movement and are less likely to move backward compared to recruits. In contrast, recruits display movement patterns more akin to a diffusive walk, with an approximately uniform sampling of the three directions. Consequently, scouts travel greater distances from the nest and spend more time exploring the arena, while recruits typically remain within a confined region, usually no farther than 25 cm from the nest. One may thus expect that the differences in the scouts’ and recruits’ exploratory behaviour lead to markedly different spatial spreading and coverage patterns.
@@
To answer quantitatively these questions we develop a discrete space-time random walk model on a bounded honeycomb lattice that faithfully represents the experimental arena. Since the honeycomb lattice is non-Bravais, to model movement and search in the experimental arena, we utilise the theory of random walks with internal states to capture the periodic heterogeneity in the domain sites marris2023exact ; montroll1969random and include a one-step memory in the model via the theory of persistent random walks marris2024persistent ; larralde2020first . We use this model first with a single walker to investigate the spreading (mean square displacement and mean number of distinct visited sites) in the infinite lattice, quantifying how quickly scouts and recruiters explore the space compared to a diffusive walker. We then study the so-called first first-passage time (fFPT) distribution of a population of independent and heterogeneous walkers. We do so by extending well-known theory weiss1983order to account for a population of distinct walkers, e.g. unequal movement statistics. We further generalise the theory to consider multiple targets and calculate the time-dependent splitting probability, defined as the probability as a function time that one member of the population reaches specific target(s) before any other individual has reached any other target. Similarly to the case of one single walker tejedor2012optimizing ; marris2024persistent , with multiple walkers one may expect that an increase in persistence hinders search times. Using the splitting probability we show in fact that a highly persistent population may miss close-by targets. In contrast, when one uses experimentally relevant parameters, we show that the local target is found first.
@@
To gain a quantitative understanding we study, via a semi-analytic procedure, search behaviour on an exact replica of the experimental arena with hard reflecting boundaries. We explore the fFPT distribution for the $\mathcal{N}$ ants observed in nine different experimental realisations, and demonstrate that in the majority of cases, the empirical measurements coincide much closer to the mode of the fFPT distribution than its mean. Moreover, by identifying the differences in the ordering of events in nest-to-patch and patch-to-patch search, we elucidate that the empirical timescales to find the second patch match closer with the nest-to-patch case.
@@
Figure 1: Subset of empirical trajectories on the Y-maze from scouts (left) and recruits (right). The purple diamond on the bottom center depicts the entry point to the experimental arena (2x1m). The purple hexagons are the locations for the food patches, changing across different experimental trials. The color gradient in the trajectories represents the temporal evolution, indicating where the trajectories start (blue) and end (yellow). Trajectories displaying an average distance to the nest exceeding 25 cm (dashed semicircle) were classified as scouts or recruits otherwise. The thickness of the zig-zag lines aids visualization of time: thick segments indicate early times, while thin segments correspond to late times).
@@ 2 Experimental protocol in the Y-maze arena
Foraging experiments were carried out with Aphaenogaster senilis, a Mediterranean ant species displaying flexible foraging strategies, ranging from individual to group recruitment Cerda2009 . To reproduce the natural spatiotemporal scales at which these ants forage, we recorded three hours of video per trial, capturing their movements within a 2x1 meters honeycomb-like arena (see Figure 1 in ref. cristin2024spatiotemporal ). This Y-maze pattern is widely utilized in studies of animal decision-making (e.g. Czaczkes2018 ), and simplifies the quantification of animal behaviour and movement since at each vertex the walker has only three choices: to go left, right or backwards. Additionally, spatial discreteness is well-suited for incorporating network theory and other statistical physics frameworks such as spin glasses Cristin2020 ; cristin2024spatiotemporal .
@@
We tracked individual ant trajectories from the videos at a temporal resolution of 2 Hz. We analysed several aspects of their movement and classified them into scouts or recruits according to the tracks’ average distance to the nest. After, the tracks were discretised and each position was assigned to the closest node (Y-shaped intersection) in the lattice. We then estimated the turning probabilities at 1-second intervals, computing whether the ant remained in the same node, or turned to the right ($r$), left ($l$), or backward ($b$), relative to its last position. This was performed separately for scouts and recruits. The relevant parameters for the recruits are $r\approx 0.0818$, $b\approx 0.0539$, $l\approx 0.0821$, and $c\approx 0.7822$ is the probability of remaining in the site, while for the scouts we found $r\approx 0.1039$, $b\approx 0.0358$ and $l\approx 0.1008$, and $c\approx 0.7595$, where all movement parameters are rounded to 4 decimal places. In Figure 1 we represent some example tracks for both scouts and recruits, to showcase their different movement patterns.
@@ 3 Master Equation for a correlated walk in a honeycomb lattice in unbounded space
To model the correlated movement of the ants on the non-Bravais honeycomb lattice we use the theory of lattice random walk with internal states montroll1969random ; weiss1994aspects . The model requires six internal states, representing both the physical site types (spatial heterogeneity) marris2023exact ; zumofen1982energy and the direction last travelled (one-step memory) marris2024persistent ; larralde2020first . For the spatial degree of freedom we follow Montroll’s procedure of embedding the honeycomb lattice into the $\mathbb{Z}^{2}$ square lattice montroll1969random , where each unit cell contains two site types ($\triangleright$ and $\triangleleft$), referred to henceforth as positive $(\mathfrak{p})$ and negative $(\mathfrak{n})$ sites. As described in Sec. 2, the movement parameters of the model represent the probability of making a right turn $(r)$ a left turn $(l)$ or a backtrack $(b)$. Since the direction previously travelled must be known to determine in which direction a these turns will take the walker, each physical site is then decomposed into three internal states representing the movement direction last travelled. We also allow the walker at each time step to remain on a site with probability $c$, such that $l+r+b+c=1$. This construction allows to express the non-Markov movement steps on the honeycomb lattice as a higher-order Markov chain van1998remarks .
@@
Equation (1) is completely general, modelling a wide range of movement behaviour. The diffusive limiting case treated in montroll1969random ; marris2023exact ; zumofen1982energy can be recovered by taking $r=l=b=q/3$ ($0<q\leq 1$). In contrast, if $r=1$ or $l=1$, the motion is deterministic whereby the walker traverses the perimeter of one hexagon in a prescribed direction for all times.
@@
where here and elsewhere capitalised bold symbols denote column vectors. We introduce the initial condition $\bm{P}(n_{1},n_{2},0)=\delta_{\bm{n},\bm{n_{0}}}\bm{U}_{m_{0}}$, which states the walkers are intitally localised at some site $\bm{n}$, with the probability of beginning their trajectory in some internal state $m$ distributed by $\bm{U}_{m_{0}}$. In our case, the elements of $\bm{U}_{m_{0}}$ are dependent on whether the walker starts in a positive or negative state as well as in which direction the walker is assumed to have entered that site from. With little information on the direction travelled at $t=0$ within the nest, we assume the walker entered via a right, left or back turn with equal probability leading to $\bm{U}_{\mathfrak{p}}=[1/3,1/3,1/3,0,0,0]^{\intercal}$ if the walker started in a positive state and $\bm{U}_{\mathfrak{n}}=[0,0,0,1/3,1/3,1/3]^{\intercal}$ otherwise.
@@
$\widehat{\widetilde{\bm{P}}}_{\bm{n_{0}},m_{0}}(\xi_{1},\xi_{2},z)=\text{e}^{i\left[\xi_{1}n_{01}+\xi_{2}n_{02}\right]}\left[\mathbb{I}-z\bm{\lambda}(\xi_{1},\xi_{2})\right]^{-1}\cdot\bm{U}_{m_{0}},$ (3)
@@
The arrows around the matrix $\bm{\lambda}(\xi_{1},\xi_{2})$ represent the movement on the honeycomb lattice encoded by the element in the matrix where $\bm{\lambda}(\xi_{1},\xi_{2})_{i,j}$ governs the movement from element $j$ to element $i$. To illustrate, consider the element $\bm{\lambda}(\xi_{1},\xi_{2})_{5,2}$. To move from state 2 to state 5, one must hop from a positive state to a negative state in the unit cell above via a backtracking from the previous movement direction. This corresponds to a movement in the South-West direction followed by a movement back in the North-East direction as may be evinced by the arrows surrounding Eq. (4) and the schematic in Fig. 2. We note that the choice of the ‘positive’ Fourier transform , $\widehat{f}(\xi)=\sum_{n\in\mathbb{Z}}\text{e}^{i\xi n}f(n)$, leads to exponents in the structure function, which reflects the direction of movement. For example, an element $p\text{e}^{i\xi_{1}}$ represents the possibility of a step to the right with probability $p$ klafter2011first . An alternative definition of the Fourier transform with negative exponent may be taken instead LucaPRX , which leads to an equivalent representation of the dynamics as Eq. (3), whereby $(\xi_{1},\xi_{2})\rightarrow(-\xi_{1},-\xi_{2})$.
@@ 4 Coverage Metrics
There are two fundamental quantities that elucidate how efficiently a random walker explores the space over time: the mean-squared displacement (MSD), $\Delta\bm{n}(t)$, and the mean number of distinct visited sites (MNDVS) up to time $t$, $S(t)$. We study here both quantities for scouts and recruits on the unbounded lattice.
@@ 4.1 Mean Squared Displacement
Some other limiting cases include the backtracking limit ($b=1$) and the circular limit ($r=1$ or $l=1$). In the former case, we find $\Delta\bm{n}(t)=3^{-1}\left[1+(-1)^{t}\right]$, which represents a walker which hops between two sites for all time. In the latter case the hexagonal structure of the honeycomb lattice leads to chiral motion, whereby the walker traverses the perimeter of one hexagon for all time. Hence, the MSD is a piecewise linear function, $\Delta\bm{n}(t)=9^{-1}\left[9-(-1)^{t}-8\cos\left(\frac{\pi t}{3}\right)\right]$ which periodically returns to zero with period six. In contrast, in the so-called directed $r=l=1/2$ case we find $\Delta\bm{n}(t)=9^{-1}\left[2^{(4-t)}-(-1)^{t}+3(-5+4t)\right]$, with super-linear dependence at short times before diffusive spread quickly takes over.

### v2

@@ Abstract
Individual behavioural heterogeneity is key to understanding how natural populations adapt flexibly to environmental changes. In ant species like Aphaenogaster senilis, two key roles in collective foraging are commonly recognized: scouts, who discover food patches, and recruits, who exploit these patches and transport food back to the nest. These roles involve distinct movement patterns and exploratory behaviours. In this chapter, we develop a correlated random walk model on a bounded honeycomb lattice to interpret and replicate empirical observations of foraging ants in an enclosed arena with honeycomb tiling. We do so by extending the theory of first-passage processes for $\mathcal{N}$ random walkers when individuals belong to a heterogeneous population. We apply this theory to examine how individual behavioural heterogeneity in ants affects collective foraging efficiency, focusing on first-passage time statistics for nest-to-patch and patch-to-patch movements. With the combined use of the mathematical model and the controlled experimental setup we evaluate (i) the impact of distinct movement strategies by scouts and recruits on finding food items and (ii) whether ants practice strict central place foraging or utilise previously discovered patches as starting points for further exploration.
@@ 1 Introduction
Social groups are rarely uniform or cohesive units; instead, individuals within these groups often exhibit consistent behavioural differences, commonly referred to as “personalities” Jolles2020 . Understanding individual heterogeneity is essential for grasping the variability in life histories and population dynamics chesson1991 . Moreover, it significantly influences the functionality of key ecological responses. In eusocial species, such as ants and termites, individual heterogeneity is exploited to associate different roles to different tasks antecologybook , e.g., typically one female and several males are reproductively active, while nonbreeding individuals care for the young, protect the colony, and provide resources. A key debate in this regard is whether these roles represent specialised task forces designed to address different needs or if individuals switch between tasks to adapt to inherent environmental uncertainties Gordon2016 .
In the context of foraging processes, ants are well known to have evolved various strategies for efficient localisation and exploitation of food resources antecologybook with two common strategies being group recruitment and mass recruitment. In group recruitment, a scout discovers a food source and returns to the nest to recruit a small number of nestmates by leaving a chemical trail that only the scout itself can trace. This strategy, where nestmates are followers, suits smaller or scattered food sources. On the other hand, mass recruitment, where the scout lays a strong pheromone trail that may attract many nestmates, is faster and involves more ants, ideal for large or abundant food sources. Key determinants to select the appropriate strategy are the environmental conditions and the type of food available. For instance, some species primarily rely on individual foraging but switch to group or mass recruitment as food sources increase in abundance and predictability, allowing ant colonies to optimise resource acquisition across different scenarios. The number of scouts and recruits involved in food collection is also flexible, varying dynamically as the food is being collected and with the amount and predictability of resources. This adaptability raises intriguing questions about the role of these distinct behaviours, how colonies adjust their numbers, and how these adjustments impact collective success.
@@
A. senilis has been selected because of its known group recruitment foraging strategies whereby scouts and recruits have distinct roles campos2014 ; martin2024 . Recent studies on a Y-maze arena have identified significant disparities in the movement patterns of these two types of ants pol2024liquid , with quantifiable differences in the probabilities of moving left, right, or backward at each step. Scouts exhibit less sinuous movement and are less likely to move backward compared to recruits. In contrast, recruits display movement patterns more akin to a diffusive walk, with an approximately uniform sampling of the three possible directions at each vertex of the honeycomb lattice. Consequently, scouts travel greater distances from the nest and spend more time exploring the arena, while recruits typically remain within a confined region, usually no farther than 25 cm from the nest. One may thus expect that the differences in numbers and exploratory behaviours of scouts and recruits would lead to markedly different spatial spreading and coverage patterns.
@@
To answer quantitatively these questions we develop a discrete space-time, non-Markovian, random walk model on a bounded honeycomb lattice that faithfully represents the experimental arena. Since the honeycomb lattice is non-Bravais, to model movement and search in the experimental arena, we utilise the theory of random walks with internal states to capture the periodic heterogeneity in the domain sites marris2023exact ; montroll1969random and include a one-step memory in the model via the theory of persistent random walks larralde2020first ; marris2024persistent . We use this model first with a single walker to investigate the spreading (mean square displacement and mean number of distinct visited sites) in the infinite lattice, quantifying how quickly scouts and recruiters explore the space compared to a diffusive walker. We then study the so-called first first-passage time (fFPT) distribution of a population of independent and heterogeneous walkers. We do so by extending well-known theory weiss1983order to account for a population of distinct walkers, that is with unequal movement statistics. We further generalise the theory to consider multiple targets and calculate the time-dependent splitting probability, defined as the probability as a function time that one member of the population reaches specific target(s) before any other individual has reached any other target. Similarly to the case of one single walker marris2024persistent ; tejedor2012optimizing , with multiple walkers one may expect that an increase in persistence hinders search times. Using the splitting probability we show in fact that a highly persistent population may miss close-by targets. In contrast, when one uses experimentally relevant parameters, we show that the local target is found first.
@@
By employing a semi-analytic procedure, we study the ant search behaviour on an exact replica of the experimental arena with hard reflecting boundaries. We explore the fFPT distribution for the $\mathcal{N}$ ants observed in nine different experimental realisations, and demonstrate that in the majority of cases, the empirical measurements coincide much closer to the mode of the fFPT distribution than its mean. Moreover, by identifying the differences in the ordering of events in nest-to-patch and patch-to-patch search, we elucidate that the empirical timescales to find the second patch match closer with the nest-to-patch case.
@@
Figure 1: Subset of empirical trajectories on the Y-maze from scouts (left) and recruits (right). The purple diamond on the bottom centre depicts the entry point to the experimental arena ($2\times 1\,\text{m}^{2}$). The purple hexagons are the locations for the food patches, changing across different experimental trials. The colour gradient in the trajectories represents the temporal evolution, indicating where the trajectories start (blue) and end (yellow). Trajectories displaying an average distance to the nest exceeding 25 cm (dashed semicircle) were classified as scouts or recruits otherwise. The thickness of the zig-zag lines aids visualization of time: thick segments indicate early times, while thin segments correspond to late times.
@@ 2 Experimental protocol in the Y-maze arena
Foraging experiments were carried out with Aphaenogaster senilis, a Mediterranean ant species displaying flexible foraging strategies, ranging from individual to group recruitment Cerda2009 . To reproduce the natural spatiotemporal scales at which these ants forage, we recorded three hours of video per trial, capturing their movements within a $2\times 1\,$ meters squared honeycomb-like arena (see Figure 1 in ref. cristin2024spatiotemporal ). This Y-maze pattern is widely utilized in studies of animal decision-making (e.g., Czaczkes2018 ), and simplifies the quantification of animal behaviour and movement since at each vertex the walker has only three choices: to go left, right or backwards. Additionally, spatial discreteness is well-suited for incorporating the present random walk theory, network theory and other statistical physics frameworks such as spin glasses Cristin2020 ; cristin2024spatiotemporal .
@@
We tracked individual ant trajectories from the videos at a temporal resolution of 2 Hz. We analysed several aspects of their movement and classified them into scouts or recruits according to the tracks’ average distance to the nest. After, the tracks were discretised and each position was assigned to the closest node (Y-shaped intersection) in the lattice. We then estimated the turning probabilities at 1-second intervals, computing whether the ant remained in the same node ($c$), or turned to the right ($r$), left ($l$), or backward ($b$), relative to its last position. This was performed separately for scouts and recruits. The relevant parameters for the recruits are $r\approx 0.0818$, $b\approx 0.0539$, $l\approx 0.0821$, and $c\approx 0.7822$, while for the scouts we found $r\approx 0.1039$, $b\approx 0.0358$ and $l\approx 0.1008$, and $c\approx 0.7595$, where all movement parameters are rounded to 4 decimal places. In Figure 1 we represent some example tracks for both scouts and recruits, to showcase their different movement patterns.
@@ 3 Master Equation for a correlated walk in a honeycomb lattice in unbounded space
To model the correlated movement of the ants on the non-Bravais honeycomb lattice we use the theory of lattice random walk with internal states montroll1969random ; weiss1994aspects . The model requires six internal states, representing both the physical site types (spatial heterogeneity) marris2023exact ; zumofen1982energy and the direction last travelled (one-step memory) larralde2020first ; marris2024persistent . For the spatial degree of freedom we follow Montroll’s procedure of embedding the honeycomb lattice into the $\mathbb{Z}^{2}$ square lattice montroll1969random , where each unit cell contains two site types ($\triangleright$ and $\triangleleft$), referred to henceforth as positive $(\mathfrak{p})$ and negative $(\mathfrak{n})$ sites. As described in Sec. 2, the movement parameters of the model represent the probability of making a right turn $(r)$ a left turn $(l)$ or a backtrack $(b)$. Since the direction previously travelled must be known to determine in which direction the walker will turn, each physical site is then decomposed into three internal states representing the movement direction last travelled. We also allow the walker at each time step to remain on a site with probability $c$, such that $l+r+b+c=1$. This construction allows to express the non-Markov movement steps on the honeycomb lattice as a higher-order Markov chain van1998remarks .
@@
Equation (1) is completely general, modelling a wide range of movement behaviour. The diffusive limiting case treated in marris2023exact ; montroll1969random ; zumofen1982energy can be recovered by taking $r=l=b=q/3$ ($0<q\leq 1$). In contrast, if $r=1$ or $l=1$, the motion is deterministic whereby the walker traverses the perimeter of one hexagon in a prescribed direction for all times.
@@
where here, and elsewhere, capitalised bold symbols denote column vectors and where the matrices containing the transition probabilities, viz. $\mathbb{A},\mathbb{B}...,\mathbb{E}$, are defined in Appendix 1. We introduce the initial condition $\bm{P}(n_{1},n_{2},0)=\delta_{\bm{n},\bm{n_{0}}}\bm{U}_{m_{0}}$, which states the walkers are intitally localised at some site $\bm{n}$, with the probability of beginning their trajectory in some internal state $m$ distributed by $\bm{U}_{m_{0}}$. In our case, the elements of $\bm{U}_{m_{0}}$ are dependent on whether the walker starts in a positive or negative state as well as in which direction the walker is assumed to have entered that site from. With little information on the direction travelled at $t=0$ within the nest, we assume the walker entered via a right, left or back turn with equal probability leading to $\bm{U}_{\mathfrak{p}}=[1/3,1/3,1/3,0,0,0]^{\intercal}$ if the walker started in a positive state and $\bm{U}_{\mathfrak{n}}=[0,0,0,1/3,1/3,1/3]^{\intercal}$ otherwise.
@@
$\widehat{\widetilde{\bm{P}}}_{\bm{n_{0}},m_{0}}(\xi_{1},\xi_{2},z)=\text{e}^{-i\left[\xi_{1}n_{01}+\xi_{2}n_{02}\right]}\left[\mathbb{I}-z\bm{\lambda}(\xi_{1},\xi_{2})\right]^{-1}\cdot\bm{U}_{m_{0}},$ (3)
@@
The arrows around the matrix $\bm{\lambda}(\xi_{1},\xi_{2})$ represent the movement on the honeycomb lattice encoded by the element in the matrix where $\bm{\lambda}(\xi_{1},\xi_{2})_{i,j}$ governs the movement from element $j$ to element $i$. To illustrate, consider the element $\bm{\lambda}(\xi_{1},\xi_{2})_{5,2}$. To move from state 2 to state 5, one must hop from a positive state to the negative state in the unit cell above via a backtracking from the previous movement direction. This corresponds to a movement in the South-West direction followed by a movement back in the North-East direction as may be evinced by the arrows surrounding Eq. (4) and the schematic in Fig. 2. We note that the choice of the ‘positive’ Fourier transform, $\widehat{f}(\xi)=\sum_{n\in\mathbb{Z}}\text{e}^{i\xi n}f(n)$, leads to exponents in the structure function, which reflects the direction of movement. For example, an element $p\text{e}^{i\xi_{1}}$ represents the possibility of a step to the right with probability $p$ klafter2011first . An alternative definition of the Fourier transform with negative exponent may be taken instead LucaPRX , which leads to an equivalent representation of the dynamics as Eq. (3), whereby $(\xi_{1},\xi_{2})\rightarrow(-\xi_{1},-\xi_{2})$.
@@ 4 Coverage Metrics
There are two fundamental quantities to elucidate how efficiently a random walker explores the space over time: the mean-squared displacement (MSD), $\Delta\bm{n}(t)$, and the mean number of distinct visited sites (MNDVS) up to time $t$, $S(t)$. We study here both quantities for scouts and recruits on the unbounded lattice.
@@ 4.1 Mean Squared Displacement
Some other limiting cases include the backtracking limit ($b=1$) and the circular limit ($r=1$ or $l=1$). In the former case, we find $\Delta\bm{n}(t)=3^{-1}\left[1+(-1)^{t}\right]$, which represents a walker which hops between two sites for all time. In the latter case, the hexagonal structure of the honeycomb lattice leads to circular-like motion, whereby the walker traverses the perimeter of one hexagon for all time. Hence, the MSD is a piecewise linear function, $\Delta\bm{n}(t)=9^{-1}\left[9-(-1)^{t}-8\cos\left(\frac{\pi t}{3}\right)\right]$ which periodically returns to zero with period six. In contrast, in the so-called directed $r=l=1/2$ case we find $\Delta\bm{n}(t)=9^{-1}\left[2^{(4-t)}-(-1)^{t}+3(4t-5)\right]$, with super-linear dependence at short times before diffusive spread quickly takes over.